Net Income and Stockholders' Equity for Four Businesses

Four different companies—Chang, Henry, Nagel, and Wilcox—show the same balance sheet at the beginning and end of a year. These data, exclusive of the amount of stockholders’ equity, are summarized as follows:

Total Assets Total Liabilities
Beginning of the year $80,300 $32,100
End of the year 128,500 54,000

On the basis of the preceding data and the following additional information for the year, determine the net income (or loss) of each company for the year.

Company Chang: No additional capital stock was issued, and no dividends were paid.
Company Henry: No additional capital stock was issued, but dividends of $5,500 were paid.
Company Nagel: Capital stock of $12,100 was issued, but no dividends were paid.
Company Wilcox: Capital stock of $15,400 was issued, and dividends of $5,800 were paid.

Select "Net Income" or "Net Loss" from the dropdown menus below. Enter all amounts as positive numbers.

Company Chang $
Company Henry $
Company Nagel $
Company Wilcox $
Answer #1

Equity at the beginning of the year for each company = $80,300 - $32,100 = $48,200

Equity at the end of the year for each company = $128,500 - $54,000 = $74,500

Calculate net income or loss for each company as follows:

Company Chang Net Income 26300
Company Henry Net income 31800
Company Nagel Net income 14200
Company Wilcox Net income 16700

Above figures have been calculated in the following manner:

Company Chang Net Income =74500-48200
Company Henry Net income =74500+5500-48200
Company Nagel Net income =74500-12100-48200
Company Wilcox Net income =74500+5800-15400-48200